OpenStudy (anonymous):

How many ways can a teacher arrange four students in the front row with a total of 30 students?

OpenStudy (anonymous):

I'm confused

OpenStudy (anonymous):

Is this a combination or a permutation? What do you think?

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

Permutation?

OpenStudy (anonymous):

Does the order in which the four students placed in the front row matter?

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

So let's say that Bill, Jim, Sue, and Betty are chosen to sit in the front row. Does it matter in what order they sit?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Great. So, as long as those four are chosen, it doesn't matter in which order they're chosen. If the order doesn't matter, is that a combination or a permutation?

OpenStudy (anonymous):

Permutation?

OpenStudy (anonymous):

Sorry, no. If the order doesn't matter, then it's a combination. If the order does matter, then it's a permutation. So this question is about calculating a combination. Do you know how to do that?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

The number of combinations of choosing r items out of a set on n items is given by\[_{n}C _{r}=\frac{ n! }{ r!\left( n-r \right)! }\]Do you understand this equation?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

OK. In this question, n is the total number of people. How many is that?

OpenStudy (anonymous):

30

OpenStudy (anonymous):

Yup. And r is the number of people that are chosen to sit in the front row. How many is that?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

Right. So the calculation becomes\[_{30}C _{4}=\frac{ 30! }{ 4!\left( 30-4 \right)! }=\frac{ 30! }{ 4!26! }\] Can you calculate this answer?

OpenStudy (anonymous):

Not really

OpenStudy (anonymous):

Do you know what the factorial symbol (!) means?

OpenStudy (anonymous):

4!x26! ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I'm using a scientific calculator

OpenStudy (anonymous):

Great. So you have\[_{30}C _{4}=\frac{ 30\times29\times28\times27\times26\times25\times24\times...\times2\times1 }{\left( 4\times3\times2\times1 \right)\left( 26\times25\times24\times...\times2\times1 \right) }\]Make sense?

OpenStudy (anonymous):

If so, there are a lot of common factors in the numerator and denominator that cancel out.

OpenStudy (anonymous):

Still there?

OpenStudy (anonymous):

Yes sorry had to do something

OpenStudy (anonymous):

Cancel out the common factors and calculate the result. What do you get?